# Ab Initio theory of Electron-phonon-coupling-induced Giant Magnetic Moments of Chiral Phonons in Magnetic Materials

### v2

@@ Abstract
Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention. However, a long-standing critical issue in this field is the lack of ab initio method to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC). Here, we resolve this challenge by developing an ab initio theory for calculating EPC-induced phonon magnetic properties, applicable to both insulating and metallic materials. Based on this theory, we demonstrate EPC-induced giant phonon magnetic moments and resulting phonon Zeeman splittings in magnetic metals, which are orders of magnitude larger than classical predictions from point-charge models. Interestingly, these splittings could open observable topological gaps in phonon spectra of magnets, generating intrinsic phonon Chern states. Through first-principles calculations, we first propose candidate materials with such intrinsic phonon Chern states hosting robust edge phonon currents which may be applied to detecting neutral particles such as dark matter particles. Our work not only establishes a theoretical foundation for EPC-induced phonon magnetic properties, but also enables the ab initio calculation of long-sought TRS-breaking phonon spectra throughout the Brillouin zone in realistic materials.
@@
Introduction. Chiral phonons, collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and Niu [1, 2] and subsequently verified experimentally in monolayer WSe2 [3]. Recently, there has been a surge of research interest in chiral phonons, covering both theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14] and experiments [15, 16, 17, 18, 19, 20, 21, 22, 23]. Of particular importance are phonon magnetic moments (PMMs) of chiral phonons and their associated phonon Zeeman splittings, playing a crucial role in chiral-phonon-involved effects [1, 24, 25, 26, 27]. However, classical models based on point-charge approximations yield extremely small PMM values due to the significant large mass of nucleus  [9]. While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [7, 9, 12], an ab initio method capable of accurately calculating EPC-induced PMMs in general materials is still lacking.
@@
Figure 1: Schematic illustrations of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-induced PMM. In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions, which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons. The classical Zeeman splitting of chiral phonons deduced from the point-charge model is negligible as a result of the significant atomic mass, whereas EPC-deiven Zeeman splitting is orders of magnitude larger. This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
@@
Besides magnetic moments, phonons can also exhibit Hall effects [28, 29, 30, 31] and topological characteristics, such as Berry curvature and Chern numbers [32, 33, 34]. When the time reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, featuring topologically protected phonon edge currents with significantly enhanced detectability compared to conventional phonon Hall effects. Nevertheless, TRS-breaking external fields for phonons, such as strong external magnetic fields or Coriolis forces, often face experimental limitations [35, 36]. Consequently, phonon Chern states driven by the phonon Zeeman splitting in magnets without external fields are highly desirable. However, predicting such intrinsic topological phonon states in magnets remains challenging due to the inability to perform ab initio calcualtions of TRS-broken phonon spectra [37, 38, 39, 40, 41, 42].
@@
To resolve these challenges, we establish an ab initio theory for determining EPC-induced phonon magnetic properties, which enables first-principles calculations of previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals. Our theory relies on treating the angular momentum of chiral phonon as an effective perturbative field to the electron-phonon coupled system, with the phonon Zeeman energy interpreted as the response to this perturbation. Based on this theory, we demonstrates EPC-induced giant PMMs leading to large phonon Zeeman splittings in magnetic metals, which are orders of magnitude stronger than classical predictions from point-charge models. Building on these splittings, we propose the inertially decoupled model to search for intrinsic phonon Chern states in realistic magnetic materials. Furthermore, through first-principles calculations, we propose promising candidate materials exhibiting intrinsic phonon Chern states, including $\mathrm{EuSi_{2}}$, $\mathrm{GdSi_{2}}$ and $\mathrm{CrB_{2}}$. Our work establishes a ab initio theoretical foundation for investigating EPC-induced phonon magnetic properties and TRS-breaking topological phonon spectra in realistic materials, which could pave the way for futher developments of chiral phonon physics.
@@
Ab initio theory for phonon Zeeman energy. Previous theories regarding the magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals and are only effective for specific high-symmetry phonon wave vectors [7, 9, 10, 12]. However, since EPC is typically much more pronounced in metallic systems, there is a pressing need to develop an ab initio theory that can comprehensively describe phonon magnetic properties in insulators, semimetals, and metals.
@@
Instead of focusing on calculating effective PMM, we choose to directly calculate experimentally observable Zeeman splittings. We begin with a TRS-breaking perturbation phonon Hamiltonian, $\delta H=-\sum_{ab}\eta_{ab}p_{a}u_{b}$, known as the Raman spin-phonon interaction [43], where $\eta_{ab}$ is a real antisymmetric matrix that satisfies $\eta_{ab}=-\eta_{ba}$. Here, $u_{a}=\sqrt{m}x_{a}$ represents the renormalized atomic displacement coordinate and $p_{a}$ represents the canonical momentum related to $u_{a}$, where $m$ denotes the atomic mass. By identifying the phonon angular momentum as $s^{\text{ph}}_{c}=\sum_{ab}\epsilon_{abc}u_{a}p_{b}$ (c is the direction of the phonon angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express the phonon Zeeman energy in a compact form as $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}$, where we have defined $\eta_{c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{ab}$. Based on the expression of phonon Zeeman energy, we propose that the angular momentum of chiral phonon can be treated as an effective perturbative field to the electron-phonon coupled system. Therefore, the Zeeman energy of chiral phonon can be regarded as an effective linear response to its angular momentum.
@@
Remarkably, $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of chiral phonons induced by EPC. Within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where the $\chi_{c}^{-}(\bm{q},\omega)$ component is defined as $\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$. The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig. 1(b)[see the supplement material (SM)] [44],
@@
$\chi_{c,s}(\bm{q},\omega)=\frac{2i\omega}{\hbar N}\sum_{\bm{k},\alpha\beta}\left|{V_{c,s}^{\bm{k},\alpha\beta}}(\bm{q},\omega)\right|^{2}\frac{f(\epsilon_{s}^{\bm{k}+\bm{q},\alpha})-f(\epsilon_{s}^{\bm{k},\beta})}{\hbar\omega-(\epsilon_{s}^{\bm{k}+\bm{q},\alpha}-\epsilon_{s}^{\bm{k},\beta})}.$ (1)
Here, $\bm{k}$ and $\bm{q}$ represent the wave vectors of the electron and phonon, respectively. $N$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic band structure with band indices $\alpha$. $f$ is the Fermi-Dirac distribution, and $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is the EPC tensor [45, 46]. Notably, all parameters in Eq. (1) can be directly derived from first-principles calculations, obviating the need for empirical inputs or additional assumptions. As a result, the previously inaccessible TRS-breaking phonon spectra driven by EPC in both insulating and metallic systems is now accessible across the whole Brillouin zone.
Phonon Zeeman splitting and enhanced phonon $g$-factor. We propose the essential conditions for realizing significant phonon Zeeman splittings by EPC. Since the primary determinant of the phonon Zeeman splitting is the susceptibility $\bm{\chi}^{-}$, there are two key points according to Eq. (1): (i) The contribution of EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is substantial, and materials composed of light atoms tend to possess strong $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$. (ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, the parametric sensitivity in Eq. (1) leads to an enhancement of $\bm{\chi}^{-}$ [44]. Consequently, we can conclude that EPC-driven large phonon Zeeman splittings are most likely to occur in narrow-gap semiconductors, semimetals and metals comprising light elements.
Large phonon Zeeman splittings suggest that phonons exhibit significant PMM. To derive the PMM, we taking the derivative $\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}$ and the phonon Zeeman energy is further rewritten as $E_{\text{Z}}=-\bm{B}\cdot\bm{\mu}^{\text{ph}}$, where $\bm{B}$ represents the effective magnetic field, and $\bm{\mu}^{\text{ph}}$ denotes the PMM (see SM  [44]),
$\mu^{\text{ph}}_{c}=-g^{\text{ph}}_{c}\frac{\mu_{B}}{\hbar}s^{\text{ph}}_{c}.$ (2)
@@
Here, $\mu_{B}$ denotes the Bohr magneton, and $g^{\text{ph}}_{c}=\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}g^{e}_{c}$ represents the effective Landé $g$-factor of the chiral phonon, where $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$, $g^{e}_{c}$ is the $g$-factor of electrons along $c$ direction involved in EPC, and the dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}$ acts as an amplification factor. This amplification factor is primarily influenced by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electronic states, and it can significantly exceed unity. This enhancement allows the $g$-factor of the chiral phonon to the order of $g^{e}_{c}$.
This EPC-enhanced PMM can be intuitively understood through the following mechanism. As chiral phonons drive atomic cyclotron motion, a portion of its angular momentum is transferred to the electrons via EPC. Owing to the considerably small mass of electrons, the angular momentum transfer gives rise to strong loop electron currents, thereby resulting in a significant PMM. In contrast, within the classical framework, the phonon’s $g$-factor is constrained to the order of $10^{-5}\sim 10^{-4}g^{e}_{c}$ due to heavy ionic masses compared to the electron mass [9, 47].
@@
Figure 2: Inertially decoupled model (IDM) and intrinsic phonon Chern states. (a) Schematic of optical chiral phonon modes: When $m\sim M$, both atoms contribute to the high-frequency model (up); in the IDM condition ($m\ll M$), the optical model is primarily contributed by light atoms (down). (b) Phonon dispersions of nonmagnetic honeycomb lattice and magnetic triangle lattice. (c-e) TRS-breaking phonon dispersions of 2D/3D IDM structures: the 2D IDM by combining the honeycomb and triangular lattices (c,d), and the 3D IDM by stacking 2D IDM along the $z$-direction (e). The color gradient from violet and yellow represents the contributions from the honeycomb and triangular lattice, and the colored gradient from blue and red represents the phonon angular momentum magnitude. (f) Edge states (cyan/magenta) of IDM under open boundaries. Parameters are detailed in the SM [44].
@@
Intrinsic phonon Chern states and the inertially decoupled model. The large phonon Zeeman splitting can pave the way toward a phonon Chern state through the phonon band inversion [44]. Following the analogy to the Chern state in the Haldane model [48], we adopt the two-dimensional (2D) honeycomb lattice [43, 49, 50] as an illustrative example. When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at $K$($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $m_{I}$ and $m_{T}$, respectively. It has been shown that when $|m_{T}|>|m_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [43].
@@
However, realizing such phonon Chern states in realistic materials hosting honeycomb lattices remains challenging. Although Coriolis forces and strong magnetic fields can produce a nonzero $m_{T}$ [43], these approaches either require large external fields or complex experimental measurements. In contrast, the EPC-enhanced phonon Zeeman splitting could offer a sizable $m_{T}$ in magnets without the need of external fields. To further satisfy the condition $|m_{T}|>|m_{I}|$, it is necessary to minimize $|m_{I}|$. Therefore, an elemental ferromagnetic honeycomb material inherently preserving inversion symmetry with $|m_{I}|=0$ would be the ideal candidate. Unfortunately, to our knowledge, no elemental ferromagnetic honeycomb material has been discovered. Furthermore, any honeycomb lattice composed of multiple elements could introduce a substantial $m_{I}$, which will make the condition $|m_{T}|>|m_{I}|$ difficult to satisfy.
@@
To seek realistic materials that exhibit intrinsic phonon Chern states, we propose a practical composite lattice model, referred to as the inertially decoupled model that comprises two types of atoms with a significant mass (inertia) difference. A prototypical example of this model is constructed by combining a honeycomb lattice composed of nonmagnetic light atoms of mass $m$ with a triangular lattice of heavy magnetic atoms of mass $M$ ($m\ll M$), as shown in Figs. 2(b,c). Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions mainly originate from the light and heavy atoms, respectively, as shown in Figs. 2(a,c), which is termed “inertially decoupled” (see SM [44] for details). Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $m_{I}$ mass at both the $K$ and $K^{\prime}$ valleys. Meanwhile, the magnetic $M$ atoms breaks TRS and provide effective magnetic fields to generate large Zeeman splittings of chiral phonons via EPC, producing the desired large $m_{T}$ mass. The large $m_{T}$ opens a topologically nontrivial phonon gap (seen in Fig. 2(c,d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by the topological edge states crossing the phonon gap, as shown in Fig. 2(f).
@@
The inertially decoupled model is also applicable to three-dimensional (3D) magnets. Interestingly, when the nonmagnetic honeycomb lattice (the $m$ sublayer) and the magnetic triangular lattice (the $M$ sublayer) are stacked alternately along the $z$-direction, each $M$ sublayer effectively acts as a buffer layer that nearly decouples the phonon dispersions of adjacent $m$ sublayers in the high-frequency regime. A representative phonon dispersion of such a 3D bulk model is shown in Fig. 2(e). Notably, since each $m$ sublayer realizes a Chern state with $|C|=1$, the system exhibits an intrinsic 3D phonon Chern state. This configuration hosts multiple chiral phonon edge states resulting from a total Chern number $|C|=N_{z}$, where $N_{z}$ is the number of $m$ sublayers.
@@
Figure 3: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$. The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$. (c,d) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (c) and with (d) the TRS breaking, respectively. (e) The phonon dispersion of $\mathrm{EuSi_{2}}$ under open-boundary conditions. (f,g) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (f) and with (g) the TRS breaking, respectively. (h) The phonon dispersion of the sandwiched slab under open-boundary conditions. The color gradient from red to blue represents the magnitude of the phonon angular momentum. Cyan and magenta label the surface or edge modes originating from opposite boundaries, respectively.
@@
Candidate materials. Building upon the our theory, we developed a dedicated computational module within Quantum ESPRESSO [51, 52] incorporating density functional perturbation theory with Hubbard corrections (DFPT+U) [45, 46] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons. Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from dynamical matrix through near-degenerate perturbation theory (see SM [44]), yielding TRS-broken phonon dispersions.

### v3

# Ab Initio Theory of Large Phonon Magnetic Moments Induced by Electron-Phonon Coupling in Magnetic Materials

@@ Abstract
Chiral phonons, characterized by nonzero angular momenta and magnetic moments, have attracted extensive attention. However, a long-standing critical issue in this field is the lack of an approach to accurately calculate phonon magnetic moments resulting from electron-phonon coupling (EPC) in realistic materials. Here, based on the linear response framework, we develop an ab initio theory for calculating EPC-induced magnetic properties of phonons, applicable to both insulating and metallic materials. Then, we evidently demonstrate EPC-induced large phonon magnetic moments and phonon Zeeman splittings resulting from spin-phonon coupling in magnetic metals. Interestingly, these splittings can open substantial topologically nontrivial phonon gaps, generating intrinsic phonon Chern states. Further, by constructing an inertially decoupled lattice model, we predict candidate materials exhibiting such intrinsic phonon Chern states with robust edge phonon currents which are proposed to detect neutral particles, such as dark matter particles. Our work enables ab initio calculations for EPC-induced magnetic properties of phonons and long-sought magnetic phonon spectra.
@@
Introduction. Chiral phonons, collective vibrations stemming from the circular motion of atoms in solids with nonzero angular momentum, were initially proposed by Zhang and Niu [1, 2] and subsequently verified experimentally in monolayer WSe2 [3]. Recently, there has been a surge of research interest in chiral phonons, covering both theoretical investigations [4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18] and experiments [19, 20, 21, 22, 23, 24, 25, 26, 27]. Of particular importance are phonon magnetic moments (PMMs) and phonon Zeeman splittings, which play a crucial role in chiral-phonon-involved effects, such as ultrafast Einstein-de Haas Effect [20] and Barnett effect [22, 28]. However, classical models based on point-charge approximations yield PMM values that are orders of magnitude below experimental measurements [11, 19]. While electron-phonon coupling (EPC) has been indicated as a promising route to enhance PMMs [8, 11, 15], an ab initio approach capable of accurately calculating EPC-induced PMMs in realistic materials is lacking.
@@
Figure 1: Schematic of the phonon Zeeman splitting driven by (a) the classical PMM and (b) the EPC-enhanced PMM. In (a), the phonon band crossing point consists of two degenerate phonon modes vibrating in two different directions (black arrows), which can be recombined into left-circularly (red) and right-circularly (blue) polarized chiral phonons. The classical phonon Zeeman splitting deduced from the point-charge model ($Z_{\text{eff}}$ is the effective charge) is negligible as a result of the significant atomic mass, whereas the EPC-induced phonon Zeeman splitting (b) is orders of magnitude larger. This significant splitting can open a substantial topologically nontrivial phonon gap, giving rise to the phonon Chern state with topological edge states [dashed lines in (b)].
@@
Besides magnetic moments, phonons can also exhibit Hall effects [29, 30, 31, 32] and topological characteristics, such as Berry curvature and Chern numbers [33, 34, 35]. When the time-reversal symmetry (TRS) is broken, phonon states with nonzero Chern numbers host a phonon analog of quantum anomalous Hall states, featuring topologically protected phonon edge currents. Nevertheless, TRS-breaking external fields for phonons, such as strong magnetic fields or Coriolis forces, often face experimental limitations [36, 37]. Consequently, phonon Chern states driven by the phonon Zeeman splitting resulting from spin-phonon coupling (SPC) in magnets without requiring external fields are highly desirable. However, it is challenging to predict such intrinsic topological phonon states in magnets due to the inability to perform ab initio calculations of TRS-breaking phonon spectra [38, 39, 40, 41, 42, 43].
@@
To resolve these challenges, we establish an ab initio theory for determining EPC-induced magnetic properties of phonons, which enables first-principles calculations of previously inaccessible TRS-breaking phonon spectra across material classes from insulators to metals. Based on this theory, we demonstrate that EPC-induced PMMs can lead to large phonon Zeeman splittings through SPC in magnetic metals, which are orders of magnitude stronger than classical predictions from point-charge models [Fig. 1(a)]. Considering that these splittings could open significant topological gaps in phonon dispersions [Fig. 1(b)], we propose an inertially decoupled lattice model to realize intrinsic phonon Chern states in realistic materials. Furthermore, through first-principles calculations, we predict a family of promising candidate materials exhibiting intrinsic phonon Chern states. Our work establishes an ab initio theoretical foundation for investigating EPC-induced magnetic properties of phonons, which paves the way for further developments of phonon physics.
@@
Ab initio theory for phonon Zeeman energy and phonon magnetic moment. Previous theories regarding magnetic properties of phonons driven by EPC have mainly been applicable to insulators or semimetals, and are restricted to the $\Gamma$ point [8, 11, 13, 15]. Since EPC is typically much more significant in metallic systems, we develop an ab initio theory that can generally calculate EPC-induced magnetic properties of phonons across the whole Brillouin zone for both metals and insulators, which is also compatible with previous theories [see the supplemental material (SM) [44]]. We will demonstrate our theory as follows.
Instead of directly calculating the magnetic moment of a single phonon, we focus on experimentally observable phonon Zeeman energy induced by the coupling between PMM and spins of electrons. We start from the total phonon Hamiltonian including the TRS-breaking perturbation as
$H=\sum_{l,a}\frac{1}{2}p_{l,a}^{2}+\frac{1}{2}\sum_{l,ab}D_{l,ab}u_{l,a}u_{l,b}-\sum_{l,ab}\eta_{l,ab}p_{l,a}u_{l,b},$ (1)
@@
where $a,b=x,y,z$ represent the Cartesian coordinates, $l=1,2,...N$ denotes the lattice site for a system containing $N$ atoms, and $u_{l,a}=\sqrt{m}x_{l,a}$ represents the renormalized atomic displacement with $m$ being the atomic mass. The last term in Eq. (1), denoted as $\delta H$, describes the TRS-breaking perturbation resulting from SPC [1], and $\eta_{l,ab}$ is a $3\times 3$ real antisymmetric matrix satisfying $\eta_{l,ab}=-\eta_{l,ba}$. In addition, $p_{l,a}=\dot{u}_{l,a}+\sum_{l,b}\eta_{l,ab}u_{l,b}$ represents the canonical momentum [44]. By identifying the phonon angular momentum as $s^{\text{ph}}_{l,c}=\sum_{ab}\epsilon_{abc}u_{l,a}p_{l,b}$ (c is the direction of angular momentum), with $\epsilon_{abc}$ being the third-order Levi-Civita tensor, we can express $\delta H$ as phonon Zeeman energy in a compact form $E_{\textrm{Z}}=\bm{\eta}\cdot\bm{s}^{\text{ph}}=\sum_{l,ab}\eta_{l,c}s^{\text{ph}}_{l,c}$, where we have defined $\eta_{l,c}=\frac{1}{2}\sum_{ab}\epsilon_{abc}\eta_{l,ab}$. Notably, we find that the angular momentum of chiral phonon can be treated as an effective perturbative field to the electron-phonon coupled system, and thus the Zeeman energy of chiral phonon can be effectively regarded as a linear response.
@@
Importantly, we will show that $\bm{\eta}$ can be derived from the TRS-breaking self-energy correction of chiral phonons. Within the framework of linear response theory, $\bm{\eta}$ is concisely expressed as $\bm{\eta}(\bm{q},\omega)=\frac{1}{i\omega}\bm{\chi}^{-}(\bm{q},\omega)$, where $\chi_{c}^{-}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)-\chi_{c,\downarrow}(\bm{q},\omega)$ [see SM [44]], indicating that the spin polarization of electrons leads to Zeeman splittings of chiral phonons. The susceptibility $\chi_{c,s}(\bm{q},\omega)$ for $s=\uparrow,\downarrow$ can be obtained through the Feynman diagram illustrated in Fig. 1(b),
@@
$\chi_{c,s}(\bm{q},\omega)=\frac{2i\omega}{\hbar N_{\bm{k}}}\sum_{\bm{k},\alpha\beta}\left|{V_{c,s}^{\bm{k},\alpha\beta}}(\bm{q},\omega)\right|^{2}\frac{f(\epsilon_{s}^{\bm{k}+\bm{q},\alpha})-f(\epsilon_{s}^{\bm{k},\beta})}{\hbar\omega-(\epsilon_{s}^{\bm{k}+\bm{q},\alpha}-\epsilon_{s}^{\bm{k},\beta})}.$ (2)
Here, $\bm{k}$ and $\bm{q}$ represent the wave vectors of the electron and phonon, respectively, and $\omega$ is the phonon frequency. $N_{\bm{k}}$ is the total number of $\bm{k}$, and $\epsilon_{s}^{\bm{k},\alpha}$ denotes the electronic band structure with band indices $\alpha$. $f$ is the Fermi-Dirac distribution, and $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is the EPC tensor [45, 46]. It should be emphasized that all parameters in Eq. (2) can be directly obtained from first-principles calculations, without empirical inputs or additional assumptions. A non-zero $\bm{\chi}^{-}$ requires spin-polarized electronic states, for example, in ferromagnets. To realize large $\bm{\chi}^{-}$, there are two key points according to Eq. (2): (i) The contribution of EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$ is substantial, and materials composed of light atoms tend to possess strong $V_{c,s}^{\bm{k},\alpha\beta}(\bm{q},\omega)$. (ii) When the energy difference $(\epsilon^{\bm{k}+\bm{q},\alpha}-\epsilon^{\bm{k},\beta})$ approaches $\hbar\omega$, it leads to an enhancement of $\bm{\chi}^{-}$. Consequently, we can conclude that EPC-induced large phonon Zeeman splittings are most likely to occur in ferromagnetic narrow-gap semiconductors, semimetals and metals containing light elements.
Then, PMM can be derived from the above phonon Zeeman energy as (see SM [44])
$\mu^{\text{ph}}_{c}=-\frac{\partial E_{\text{Z}}}{\partial B_{c}}=-g^{\text{ph}}_{c}\frac{\mu_{B}}{\hbar}s^{\text{ph}}_{c}.$ (3)
@@
Here, $\mu_{B}$ denotes the Bohr magneton of electron, $B_{c}$ is the effective magnetic field, and $g^{\text{ph}}_{c}=\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}g^{e}_{c}$ represents the effective Landé $g$-factor of the chiral phonon (see SM [44]), where $E_{F}$ is the Fermi energy and $\chi_{c}^{+}(\bm{q},\omega)=\chi_{c,\uparrow}(\bm{q},\omega)+\chi_{c,\downarrow}(\bm{q},\omega)$. $g^{e}_{c}$ is the $g$-factor along $c$ direction of electrons involved in EPC. Equation (3) enables first-principles calculations of PMM. The dimensionless parameter $\frac{\hbar}{2i\omega}\frac{\partial\chi_{c}^{+}}{\partial E_{F}}$ acts as an amplification factor dominated by the EPC tensor $V_{c,s}^{\bm{k},\alpha\beta}$ and density of electron states near the Fermi energy [Eq. (S22) in SM [44]], and it allows the $g$-factor of the chiral phonon to reach the order of $g^{e}_{c}$, leading to a large PMM. This enhancement of PMM originates from electron loop currents induced by the EPC-mediated angular momentum transfer from ions to electrons, where electron loop currents are significantly enhanced owing to the tiny electron mass. In contrast, within the classical framework, the phonon’s $g$-factor is constrained to $10^{-5}\sim 10^{-4}$ order of $g^{e}_{c}$ due to large ionic masses compared to the electron mass [11, 47].
@@
Figure 2: Inertially decoupled model (IDM) and intrinsic phonon Chern states. (a) Schematic of chiral phonon modes in the high frequency regime: When $m\sim M$, both atoms contribute to high-frequency modes (upper panel); in the IDM condition ($m\ll M$), high-frequency modes are primarily contributed by light atoms (lower panel). (b) Phonon dispersions of nonmagnetic honeycomb lattice and ferromagnetic triangular lattice. (c-e) TRS-breaking phonon dispersions of the 2D IDM by combining the honeycomb and triangular lattices (c,d) and the 3D IDM by stacking 2D IDM along the $z$-direction (e). (f) Edge states (cyan/magenta) of the 2D IDM. Parameters are detailed in the SM [44].
@@
Intrinsic phonon Chern states and the inertially decoupled model. The accurate calculation of phonon Zeeman splitting enables the determination of magnetic phonon spectra throughout the Brillouin zone, which is indispensable for exploring magnetic phonon topology. Following the analogy to the Chern state in the Haldane model [48], we adopt the two-dimensional (2D) honeycomb lattice [49, 50, 51] as an illustrative example. When either inversion symmetry ($I$) or TRS ($T$) is broken, each phonon Dirac cone at $K$($K^{\prime}$) valley acquires a finite gap, effectively gaining masses denoted as $\Delta_{I}$ and $\Delta_{T}$, respectively. It has been shown that when $|\Delta_{T}|>|\Delta_{I}|$, band inversion occurs in one of the two Dirac cones, resulting in a topological phonon state with a Chern number of $|C|=1$ [49].
@@
However, it is challenging to realize such phonon Chern states in realistic materials with honeycomb lattices. Although Coriolis forces and strong magnetic fields can produce a nonzero $\Delta_{T}$ [49], these approaches either require large external fields or complex experimental measurements. In contrast, the phonon Zeeman splitting could offer a sizable $\Delta_{T}$ in magnets without the need of external fields. To further satisfy the condition of $|\Delta_{T}|>|\Delta_{I}|$, it is necessary to minimize $|\Delta_{I}|$. Therefore, an elemental ferromagnetic honeycomb material inherently preserving inversion symmetry with $|\Delta_{I}|=0$ would be the ideal candidate. Unfortunately, to our knowledge, no elemental ferromagnetic honeycomb material has been discovered. Furthermore, any honeycomb lattice composed of multiple elements could introduce a substantial $\Delta_{I}$, which will make the condition $|\Delta_{T}|>|\Delta_{I}|$ difficult to satisfy.
@@
To seek realistic materials that exhibit intrinsic phonon Chern states, we propose a practical composite lattice model, referred to as the inertially decoupled model that comprises two types of atoms with a significant mass (inertia) difference. A prototypical example of this model is constructed by combining a honeycomb lattice composed of nonmagnetic light atoms of mass $m$ with a triangular lattice of heavy magnetic atoms of mass $M$ ($m\ll M$), as shown in Figs. 2(b,c). Due to the difference in atomic inertia, the high-frequency and low-frequency dispersions are mainly contributed by the light and heavy atoms, respectively, as shown in Figs. 2(a,c), which is termed “inertially decoupled” (see SM [44] for details). Within the honeycomb sublattice, the inversion symmetry is preserved, resulting in a vanishing $\Delta_{I}$ mass at both the $K$ and $K^{\prime}$ valleys. Meanwhile, the magnetic $M$ atoms break TRS to generate large Zeeman splittings of chiral phonons via SPC, producing the desired large $\Delta_{T}$ mass. The large $\Delta_{T}$ opens a topologically nontrivial phonon gap (seen in Fig. 2(d)), thereby realizing an intrinsic phonon Chern state with $|C|=1$, characterized by topologically protected edge states crossing the phonon gap, as shown in Fig. 2(f).
@@
The inertially decoupled model is also applicable to three-dimensional (3D) magnets. Interestingly, when the nonmagnetic honeycomb lattice (the $m$ sublayer) and the ferromagnetic triangular lattice (the $M$ sublayer) are stacked alternately along the $z$-direction, each $M$ sublayer effectively acts as a buffer layer that nearly decouples the phonon dispersions of adjacent $m$ sublayers in the high-frequency regime. A representative phonon dispersion of such a 3D bulk model is shown in Fig. 2(e). Notably, since each $m$ sublayer realizes a Chern state with $|C|=1$, the system exhibits a 3D phonon Chern state. This configuration hosts multiple unidirectional phonon edge states resulting from a total Chern number $|C|=N_{z}$, where $N_{z}$ is the number of $m$ sublayers.
@@
Figure 3: Lattice structure and TRS-breaking phonon dispersions of ferromagnetic $\mathrm{EuSi_{2}}$. (a,b) The front view (a) and the side view (b) of the lattice structure of $\mathrm{EuSi_{2}}$. (c,d) The phonon dispersions of bulk $\mathrm{EuSi_{2}}$ without (c) and with (d) the SPC, respectively. (e) The phonon dispersion of bulk $\mathrm{EuSi_{2}}$ under open-boundary conditions. (f,g) The phonon dispersion of the sandwiched slab enclosed by the black rectangle in (b) without (f) and with (g) the SPC, respectively. (h) The phonon dispersion of the sandwiched slab under open-boundary conditions. The color gradient from red to blue represents the phonon angular momentum. Cyan and magenta colors in (c) and (h) label topological boundary modes residing on opposite boundaries.
@@
Candidate materials. Building upon our theory, we developed a dedicated computational module within Quantum ESPRESSO [52, 53] incorporating density-functional perturbation theory with Hubbard corrections (DFPT+U) [45, 46] for precise determination of response coefficients $\bm{\eta}$ and Landé $g$-factors of phonons. Then we add the effect of $\bm{\eta}$ into Schrödinger-like phonon equations constructed from the dynamical matrix through near-degenerate perturbation theory (see SM [44] for details), yielding TRS-breaking phonon dispersions.